Solve


x ^ 2 * (d ^ 2 * y)/(dx) + x * (dy)/(dx) - 9y = 48x ^ 5

1
Expert's answer
2021-12-28T11:09:58-0500

"x^2\frac{d^2y}{dx^2}+x\frac{dy}{dx}-9y=48x^5"

"D(D-1)y+Dy-9y=48x^5"

"(D^2-D+D-9)y=48x^5"

"(D^2-9)y=48x^5"

A.E;

"m^2-9=0\implies m=3,-3"

Put "x=e^z\therefore z=ln\ x"

C.F="C_1e^{-3z}+C_2e^{3z}"

"P.I=\frac{1}{(D^2-9)}\cdot48e^{5z}"

"=\frac{1*48}{((5)^2-9)}\cdot e^{5z}=\frac{48}{16}\cdot e^{5z}"

"P.I=3e^{5z}"

Complete solution;

y=C.F+P.I

"y=C_1e^{-3z}+C_2e^{3z}+3e^{5z}"

"y=C_1e^{-3ln\ x}+C_2e^{3ln\ x}+3e^{5ln\ x}"

"y=\frac{C_1}{x^3}+C_2x^3+3x^{5}" , where C1 and C2 are arbitrary constants